Prove that:
1 + cos A/ 1 - cos A = tanČA/(secA - 1)Č

Answer by kapilsinghi(68)

1 + cos A/ 1 - cos A = tanČA/(secA - 1)Č
LHS = (1 + cos A)/ (1 - cos A)
RHS = tanČA/(secA - 1)Č
= [tanA/(sec-1)]^2
= [tanA/((1/cosA)-1)]^2
= [tanA/(1-cosA)/cosA]^2
= [(tanA*cosA)/(1-cosA)]^2
= [sinA/(1-cosA)]^2
= sinA^2/(1-cosA)^2
= (1-cosA^2)/(1-cosA)^2
= [(1-cosA)*(1+cosA)]/[(1-cosA)*(1-cosA)]
= (1+cosA)/(1-cosA)
LHS = RHS
Regards
Kapil
kapilsinghi123@gmail.com